OpenStudy (anonymous):

Solve the following system of equations and show all work. y = −x^2 + 4 y = 2x + 1 Please Help

OpenStudy (anonymous):

@Hero @Here_to_Help15

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@jane11509

OpenStudy (anonymous):

@BlossomCake @butterflydreamer @Lady.Liv1776 @mathmate

OpenStudy (anonymous):

@Hero please

OpenStudy (anonymous):

-x^2+4=2x+1

OpenStudy (anonymous):

now solve for x

OpenStudy (anonymous):

Ok hold on

OpenStudy (anonymous):

I am confused

OpenStudy (anonymous):

what are you stuck on?

OpenStudy (anonymous):

Do I got to square root first or subtract 4 on both sides

OpenStudy (anonymous):

this problem involves factoring the equation

OpenStudy (anonymous):

so first get everything on one side of the equation

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

-x^2+4=2x+1

OpenStudy (anonymous):

x^2 + 2x - 3 = 0

OpenStudy (anonymous):

(x+3)(x-1)=0

OpenStudy (anonymous):

x is -3 and 1

OpenStudy (anonymous):

plug these values into one of the equations to find the corresponding y-value

OpenStudy (anonymous):

y=2(-3)+1

OpenStudy (anonymous):

y=-5

OpenStudy (anonymous):

first solution (-3,-5)

OpenStudy (anonymous):

y=2(1)+1

OpenStudy (anonymous):

y=3

OpenStudy (anonymous):

second solution (1,3)

OpenStudy (anonymous):

I'm sorry I had stepped away but thank you so much

OpenStudy (anonymous):

np